OpenStudy (anonymous):

use the value of the trig func to evaluate the indicated functions. sin (-t)= 3/8 (a) sin t (b) csc t

OpenStudy (anonymous):

\[\huge \color{green} {\sin(-x) = -sinx}\]

OpenStudy (anonymous):

then what's sin t? :O

OpenStudy (anonymous):

\[\huge \color{blue}{cosecx = \frac{1}{sinx}}\]

OpenStudy (anonymous):

Just replace x with t and use the formula..

Parth (parthkohli):

@dorkkk Do you need to understand anything more? Can I clear the mess for you?

Parth (parthkohli):

\(\Rightarrow \sin (-t) = \Large{3 \over 8} \) \(\Rightarrow -\sin(t) = \Large{3 \over 8}\) \(\Rightarrow -(-\sin (t)) = \Large{-3 \over 8}\) \(\Rightarrow\sin t = -\Large {3 \over 8} \)

Parth (parthkohli):

@dorkkk Did you get it?

OpenStudy (anonymous):

okay i looked at the table

OpenStudy (anonymous):

what about csc t?

Parth (parthkohli):

\(\csc \theta \Longrightarrow \text{Just flip the } \sin \theta \)

Parth (parthkohli):

So if \(\sin t= -\Large{3 \over 8}\), then \(\csc t = -\Large{8 \over 3}\)

Parth (parthkohli):

@dorkkk Get it? :)

OpenStudy (anonymous):

ahh, ok. thank you :D

OpenStudy (anonymous):

Thanks for solving the entire question on your own @ParthKohli ..

Parth (parthkohli):

Is that sarcasm?

Parth (parthkohli):

Oh, I apologize for solving this question. I was just explaining it to him.

OpenStudy (anonymous):

Explaining?? By solving the entire??

OpenStudy (anonymous):

If you really want to explain then you can take the other examples too..

Parth (parthkohli):

Yes. I explained him and gave the answer because it was the first time he encountered such a question. I'd ask him next time if there is a similar question related to this.

OpenStudy (anonymous):

Did he reply any of your questions earlier.?? When you asked him after solving the entire then he said thanks... Did not you get that he only wanted the answer and not the solution??

Parth (parthkohli):

Yes, I did. I'd be asking him next time, as I told you. Please don't start a fight here. Thanks.

OpenStudy (anonymous):

You can take other examples too if you want to explain... then you may say him to do the original on your own..

OpenStudy (anonymous):

woah woah woah, i don;t want any arguments here and yes, it was the first time i encountered this question. i am a girl btw :P

Parth (parthkohli):

No, I sensed that Dork is a quick learner. She answered something in the last question.

OpenStudy (anonymous):

If you are girl, then it does not mean that you will not reply if someone asks you...

Parth (parthkohli):

I know the feeling of how you are confused. Are all your doubts clear now, Dork?

OpenStudy (anonymous):

im just clearing it up because pathkohli was referring to me as a him.

Parth (parthkohli):

I have to admit — That name makes me crack every time I see it ;)

OpenStudy (anonymous):

yes, im good now and done with precal hw! chapter aka trig functions is quite a difficult chapter

OpenStudy (anonymous):

If her doubts are cleared then give her a new problem to solve @ParthKohli ..

OpenStudy (anonymous):

chapter 4*

Parth (parthkohli):

Dork, can you calculate what \(\csc 45 \) is?

OpenStudy (anonymous):

And without using a calcuator @dorkkk ..

Parth (parthkohli):

You may ask us how to do it, if you have any doubt.

OpenStudy (anonymous):

Yes sure if you have doubt then let us know..@dorkkk

OpenStudy (anonymous):

1/sin45? =√2

Parth (parthkohli):

Yes. Correct.

OpenStudy (anonymous):

cool! :) thank you for your help. good night,

Parth (parthkohli):

You're welcome! Good night!